OpenStudy (anonymous):

3x squared - 3y squared

OpenStudy (anonymous):

impossible to answer unless you give us some hints

OpenStudy (solomonzelman):

\[3x^2-3y^2~~~~~~~~~~~~~or~~~~~~~~~~~~~~~~(3x)^2-(3y)^2\]

OpenStudy (solomonzelman):

if it's the first one, factor out of 3, and use a^2-b^2=(a+b)(a-b) if it's the second one, then just use a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

You welcome!